OpenStudy (anonymous):

how do you solve this equation (20-r)^1/2=r

OpenStudy (anonymous):

naynay, you need to square both sides so you can get access to the r under the square root. So\[20-r=r^2 \rightarrow r^2+r-20=0\]which is quadratic. This can now be factored as\[(r+5)(r-4)=0\]so \[r=-5\]or\[r=4\]

OpenStudy (anonymous):

how do you solve this (6b)^1/2=(8-2b)^1/2

OpenStudy (anonymous):

One sec.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You have to remove the square roots. You do this by squaring. Whatever you do to one side, you must do to the other, so,\[\sqrt{6b}=\sqrt{8-2b} \rightarrow 6b=8-2b \rightarrow 8b=8 \rightarrow b=1\]

OpenStudy (anonymous):

ok so b will equal 1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh ok i see what you did

OpenStudy (anonymous):

ok this one i really dont get -3=(37-3n)^1/2-n

OpenStudy (anonymous):

Have you become a fan yet?

OpenStudy (anonymous):

no i will

OpenStudy (anonymous):

You're always looking to isolate your variable. Here I would add n to both sides and then square both sides, so\[n-3=\sqrt{37-3n} \rightarrow (n-3)^2=37-3n\]Then expand the left hand side\[n^2-6n+9=37-3n \rightarrow n^2-3n-28=0\]which you can then solve by factoring.

OpenStudy (anonymous):

where did you get 37 from

OpenStudy (anonymous):

You gave it to me in your question: -3=(37-3n)^1/2-n

OpenStudy (anonymous):

oh ok im sorry didnt look twice =)

OpenStudy (anonymous):

np

OpenStudy (anonymous):

thank you a lot with helping me because im really not all that good in math

OpenStudy (anonymous):

You're welcome. All you need is practice.

OpenStudy (anonymous):

Things will start to click.

OpenStudy (anonymous):

yeah i have a friend of mine comin to help me twice a week so hopefully it work =)

OpenStudy (anonymous):

I'm sure it will. Good luck, and you can use this site too for help.

OpenStudy (anonymous):

yeah i like this site a lot it does help

OpenStudy (anonymous):

can you help me with this (-3-4x)^1/2-(-2-2x)^1/2=1

OpenStudy (anonymous):

\[\sqrt{-3-4x}-\sqrt{-2-2x}=1\]Okay, for ones like this, it's best to move one of the square roots to the other side and then square both sides:\[\sqrt{-3-4x}=1+\sqrt{-2-2x}\]square both sides:\[-3-4x=\left( 1+\sqrt{-2-2x} \right)^2\]Now the hard part is to expand the right-hand side\[\left( 1+\sqrt{-2-2x} \right)^2=1+2\sqrt{-2-2x}+\left( \sqrt{-2-2x} \right)^2\]\[=1+2\sqrt{-2-2x}+(-2-2x)\]\[=-1-2x+\sqrt{-2-2x}\]

OpenStudy (anonymous):

ok so this girl helped me and we got it right =)

OpenStudy (anonymous):

Okay, so now looking at this, we have the problem of the square root again. This is because of the expansion we had to do earlier BUT we can get rid of it using the same technique as before - move everything that has nothing to do with the square root to the other side, and then square both sides.\[-3-4x=-1-2x-\sqrt{-2-2x}\]becomes\[-2-2x=-\sqrt{-2-2x}\]which is the same as\[\sqrt{-2-2x}=2+2x \rightarrow -2-2x=(2+2x)^2\]\[-2-2x=4+8x+4x^2 \rightarrow 4x^2+10x+6=0\]i.e.\[2x^2+5x+3=0 \rightarrow (x+2)(x+3)=0 \rightarrow x=-2,-3\]

OpenStudy (anonymous):

Well, I just thought I'd finish it.